# Accelerating Mobile Language Model Generation via Hybrid Context and Hardware Coordination

### v1

@@ Abstract.
Enhancing on-device large language models (LLMs) with contextual information from local data enables personalized and task-aware generation, powering use cases such as intelligent assistants and UI agents. While recent developments in neural processors have substantially improved the efficiency of prefill on mobile devices, the token-by-token generation process still suffers from high latency and limited hardware utilization due to its inherently memory-bound characteristics. This work presents CoordGen, a mobile inference framework that integrates speculative decoding with dynamic hardware scheduling to accelerate context-aware text generation on mobile devices. The framework introduces three synergistic components: (1) adaptive execution scheduling, which dynamically balances compute graphs between prefill and decoding phases; (2) context-aligned drafting, which improves speculative efficiency through lightweight online calibration to current tasks; and (3) hardware-efficient draft extension, which reuses and expands intermediate sequences to improve processing parallelism and reduce verification cost. Experiments on multiple smartphones and representative workloads show consistent improvements of up to 3.8× in generation speed and 4.7× in energy efficiency compared with existing mobile inference solutions. Component-level analysis further validates the contribution of each optimization.
@@ 1. Introduction
Figure 1. Workflow of context-augmented generation and comparison between CoordGen and related work :
@@
In this paper, we present CoordGen, the first system to enable retrieval-based speculative decoding for efficient on-device CAG through end-to-end on-chip NPU offloading. During prefill, CoordGen employs chunked prefill-based NPU inference like existing approaches (Xu et al., 2025b; Chen et al., 2025a); during decoding, it dynamically switches to a compute graph optimized for decoding, performing retrieval-driven drafting and parallel verification. Designing CoordGen requires addressing the following three unique challenges of aligning retrieval-based speculative decoding with NPUs, which prior work has not explored:
@@
$\bullet$ Progressive Graph Scheduling (Section 3.2). CoordGen deploys two specialized graphs and switches between them asynchronously to minimize overhead. The LLM is partitioned into blocks. During each round of chunked prefill, we switch to a decoding-optimized block graph while prefill computation continues in parallel. The switched decoding graph is invoked multiple times within the prefill stage to ensure result correctness. This design rests on two observations: (i) executing short-input graphs multiple times is equivalent to a single long-input execution, and (ii) graph loading can be overlapped with one-round chunked prefill computation. Consequently, CoordGen can transition to the decoding graph without additional load overhead, yielding a negligible graph-switching cost.
$\bullet$ In-Context Distribution Calibration (Section 3.3). Motivated by the observation that, during prefill, the model computes next-token logits for each context token that already align with its output distribution, CoordGen uses these logits to calibrate the task context by constructing a model-distributed token tree via depth-first search. In the subsequent retrieval-based drafting stage, CoordGen retrieves from this calibrated tree to mitigate lexical divergence between the context and the model’s expectations. Compared with costly precomputation or task-specific fine-tuning, this lightweight procedure incurs negligible overhead.
$\bullet$ NPU-Optimized Draft Reuse (Section 3.4). CoordGen extends draft construction by selectively reusing plausible tokens from previously rejected drafts based on a confidence-based strategy. This is motivated by the observation that rejected tokens may still capture correct semantics and be accepted in later decoding steps. By reusing such tokens, CoordGen increases effective draft length with useful content, improving NPU utilization and reducing verification costs.
@@
$\bullet$ We design and implement CoordGen, an efficient NPU-optimized on-device context-augmented generation framework with three novel techniques, bridging the gap between retrieval-based speculative decoding and NPU offloading.
$\bullet$ We conduct comprehensive evaluations across four datasets, three smartphones, and three LLMs, demonstrating that CoordGen delivers consistent and significant performance gains in diverse mobile scenarios.
@@ 3.1. Overview
Figure 7. Overview of CoordGen framework.:
Design goal. CoordGen accelerates context-augmented generation by NPU-optimized retrieval-based speculative decoding. This entails (1) deploying and scheduling two compute graphs optimized for prefill and decoding, (2) aligning task context with LLM distribution, and (3) enriching drafts with sufficient high-value tokens to fully utilize NPU resources.
Workflow. Figure 7 illustrates the system architecture and workflow of CoordGen.
  • Offline stage. CoordGen partitions the LLM into multiple blocks, each containing several transformer layers. It profiles the load and compute latency of each block through warm-up to determine scheduling strategy.
  • Online stage. Upon receiving a request, CoordGen fetches task-relevant database and performs prefill using a SOTA chunk-based approach (Xu et al., 2025b; Chen et al., 2025a), splitting the input into chunks processed by a long-sequence optimized graph. During prefill, progressive graph scheduling (Section 3.2) incrementally switches the prefill-optimized graph to the decoding-optimized graph with load overlapped by computation. After prefill, output logits are used for in-context distribution calibration (Section 3.3), augmenting the database with calibrated contexts. CoordGen then performs retrieval-based speculative decoding using the decoding-optimized graph and the calibrated database. It iteratively constructs a drafting tree and verifies it via a confidence-based strategy to identify and reuse valuable tokens (Section 3.4). After an end-of-sequence token is predicted, the final response is returned and stored in the database manager as a historical context for future requests.
Key techniques. CoordGen incorporates three techniques to align retrieval-based speculative decoding with NPU offloading:
@@ 3.2. Progressive Graph Scheduling
Switching graph progressively. CoordGen gradually switches $G^{1}$ to $G^{2}$ during prefill. The design builds on two insights:
  • Graph equivalence by repetition: executing $G^{2}$ multiple times on shorter inputs yields the same result as one execution of $G^{1}$ on longer inputs, allowing early switching. As CoordGen maximizes draft lengths (Section 3.4), $G^{2}$ will not be too short (e.g., chunk/2 or chunk/4 of $G^{1}$), which will not incur much performance degradation.
@@
Figure 8 illustrates progressive scheduling with $N=6$ and two input chunks. We suppose that CoordGen decides to overlap loading a block by two blocks of computation based on the profiling results. Initially, all blocks are loaded with $G^{1}$. CoordGen computes $\operatorname{P}_{1},\operatorname{P}_{2}$ with $G^{1}$ while loading $G^{2}$ for $\operatorname{P}_{6}$. Then $\operatorname{P}_{3},\operatorname{P}_{4}$ execute while $G^{2}$ loads for $\operatorname{P}_{5}$. To compute $\operatorname{P}_{5},\operatorname{P}_{6}$ under $G^{2}$, chunk 1 is split into sub-chunks matching $G^{2}$’s shape, with $G^{2}$ executed repeatedly. Prefill of chunk 1 finishes with output logits while half of the model is switched. The same procedure applies to chunk 2. In contrast, synchronous loading waits for all compute to finish, and naive asynchronous loading only overlaps in chunk 2, both incurring higher latency (Figure 8(b)).
Problem Formulation. Formally, given $N$ model blocks that sequentially execute prefill for $C$ chunks, CoordGen aims to determine the time to load each block, i.e., choosing blocks $\{\operatorname{P_{i},\dots}\}$ to overlap load overhead of $\operatorname{P_{j}}$ by their computations at each compute-load round. The objective is to minimize the overall latency, which is equal to the maximum of the prefill completion time and the switch completion time.
Scheduling methodology. This scheduling problem is NP-hard as it is essentially equivalent to the 2-Partition problem (Korf, 1998) (proofed in Appendix A). CoordGen addresses this via a greedy algorithm, choosing the least $k$ consecutive blocks $\{\operatorname{P}_{i},\dots,\operatorname{P}_{i+k-1}\}$ that overlap a block load to execute. This strategy enables fully masking the load overhead by computation to minimize switching overhead. CoordGen finds the smallest $k$ by estimating $\widehat{T}_{\text{load}}(\operatorname{P}_{j};G^{2})$ and $\widehat{T}_{\text{compute}}(\operatorname{P}_{i};G^{x}_{i})$ based on the profiling results. To hide load latency, the sum of the next $k$ blocks’ compute time must be larger than the load overhead of the last block not switched, formulated as:
@@
CoordGen greedily chooses the minimal feasible $k$ that satisfies Equation 1:
@@
Then CoordGen executes $\operatorname{P}_{i},\dots,\operatorname{P}_{i+k^{\star}-1}$ under current graphs while loading $G^{2}$ for $\operatorname{P}_{j}$. This process repeats until all blocks are switched.
@@ 3.3. In-Context Distribution Calibration
Workflow. After prefill, CoordGen first collects predicted logits for all positions. Then CoordGen builds a calibrated token tree representing the model distribution by depth-first search, which includes two steps:
@@ 3.4. NPU-Optimized Draft Reuse
Reuse plausible drafts. As highlighted in Section 2.1, the model output usually shares similar semantics with the task context despite existing lexical gaps. Although these gaps cause rejection due to token-level misalignment, the rejected tokens include correct semantics and may be accepted in the next verification. CoordGen identifies and reuses these plausible tokens to construct NPU-friendly drafts lengthened by task-relevant information.
@@
Reuse methodology. To leverage reusable tokens in the first two cases, CoordGen aims to: (1) determine the start and end positions $\gamma,\epsilon$ of the maximum reusable segment; (2) estimate the reuse lifetime $\delta$, i.e., how many iterations the reused segment should be reserved for verification.
Enumerating all possible reuse strategies is of $O(N^{3})$ complexity, which is not suitable for online scheduling. Therefore, CoordGen adopts a confidence-based and length-first reuse strategy, assuming that drafts aligned with model predictions are more likely to be accepted in future steps. Therefore, we identify and retain the longest segment of $\mathbf{x}$ such that $x_{i}=y_{i}$ for all $i\in[e+\gamma,e+\epsilon]$, by solving:
@@
By maximizing draft length with potentially acceptable tokens, this reuse mechanism directly enhances NPU utilization and reduces verification times. The overhead of CoordGen’s drafting and reusing processes is near ignorable with ¡3 ms per decoding step, as discussed in Section 4.4.
@@ 4.1. Implementation and Setup
Implementation. We implement CoordGen  on top of the mllm framework (Xu et al., 2025b) with approximately 6,000 lines of C/C++ code. We reuse the NPU backend of mllm, including neural kernels and memory managers, to perform prefill and draft verification. CoordGen inherits kernel-level and layer-level optimizations from mllm and further extends the system with speculative decoding capabilities. Our modular design allows easy porting to other mobile LLM inference frameworks that face NPU inefficiencies during decoding. To enable NPU-optimized speculative decoding, we augment mllm with customized kernels and sampling strategies. In order to accelerate drafting, we implement a suffix automaton to identify potential suffix matches (Mohri et al., 2009). To support NPU graph switching, we integrate the graph saving and loading APIs provided by the Qualcomm QNN SDK (Qualcomm, 2025a).
Hardware. We evaluate CoordGen on three representative smartphones with Qualcomm SoCs: Redmi K60 Pro (Snapdragon 8 Gen 2, Android 13, 12 GB RAM), Redmi K70 Pro (Snapdragon 8 Gen 3, Android 14, 16 GB RAM), and OnePlus 13 (Snapdragon 8 Gen Elite, Android 14, 24 GB RAM). All devices are evaluated under real mobile conditions with the CPU frequency governed by the Android OS’s dynamic voltage and frequency scaling controller.
@@
Baselines. We compare CoordGen against six baselines, all implemented atop mllm to eliminate performance gaps from backend design differences  [^1] Other existing mobile LLM inference frameworks with NPU offloading support such as PowerInfer-v2 (Xue et al., 2024) and HeteroLLM (Chen et al., 2025a) are not open-source, making them unsuitable for direct implementation. CoordGen, as a decoding optimization, is orthogonal and integrable with PowerInfer-v2 and HeteroLLM, so we do not compare it directly. EdgeLLM (Xu et al., 2025a) optimizes model-based SD mainly to alleviate memory constraints. With sufficient memory, its performance reduces to that of EAGLE and does not address the memory-bound issue of decoding, thus making it an uninformative baseline for our settings. . All baselines use NPU for prefill, which is a widely-adopted approach for both industry and academic (Xu et al., 2025b; Chen et al., 2025a; Li et al., 2024a; Xue et al., 2024). We categorize these baselines by whether they use CPU for decoding or apply speculative decoding:
  • (1)NPU vanilla: use NPU for both prefill and decoding222mllm (Xu et al., 2025b) and other on-device LLM engines  (Chen et al., 2025a; Xue et al., 2024) offload precision-sensitive operations, such as attention, to CPU/GPU to improve response quality. CoordGen follows this pattern even in the NPU vanilla mode.;
@@
Metrics and Configuration. We evaluate average per-token latency, energy consumption and peak memory. We exclude accuracy comparison as CoordGen achieves lossless generation (Leviathan et al., 2023a). Energy is measured via Android’s virtual file system under /sys/class/power_supply by profiling every 100ms. The Redmi K70 Pro is excluded from energy results as the device lacks root access. Experiments are repeated three times and we report the average numbers.

### v3

# Accelerating Mobile Language Model via Speculative Decoding and NPU-Coordinated Execution

@@ Abstract.
Enhancing on-device large language models (LLMs) with contextual information from local data enables personalized and task-aware generation, powering use cases such as intelligent assistants and UI agents. While recent developments in neural processors have substantially improved the efficiency of prefill on mobile devices, the token-by-token generation process still suffers from high latency and limited hardware utilization due to its inherently memory-bound characteristics. This work presents sd.npu, a mobile inference framework that integrates speculative decoding with dynamic hardware scheduling to accelerate context-aware text generation on mobile devices. The framework introduces three synergistic components: (1) adaptive execution scheduling, which dynamically balances compute graphs between prefill and decoding phases; (2) context-aligned drafting, which improves speculative efficiency through lightweight online calibration to current tasks; and (3) hardware-efficient draft extension, which reuses and expands intermediate sequences to improve processing parallelism and reduce verification cost. Experiments on multiple smartphones and representative workloads show consistent improvements of up to 3.8× in generation speed and 4.7× in energy efficiency compared with existing mobile inference solutions. Component-level analysis further validates the contribution of each optimization.
@@ 1. Introduction
Figure 1. Workflow of context-augmented generation and comparison between sd.npu and related work :
@@
In this paper, we present sd.npu, the first system to enable retrieval-based speculative decoding for efficient on-device CAG through end-to-end on-chip NPU offloading. During prefill, sd.npu employs chunked prefill-based NPU inference like existing approaches (Xu et al., 2025b; Chen et al., 2025a); during decoding, it dynamically switches to a compute graph optimized for decoding, performing retrieval-driven drafting and parallel verification. Designing sd.npu requires addressing the following three unique challenges of aligning retrieval-based speculative decoding with NPUs, which prior work has not explored:
@@
$\bullet$ Progressive Graph Scheduling (Section 3.2). sd.npu deploys two specialized graphs and switches between them asynchronously to minimize overhead. The LLM is partitioned into blocks. During each round of chunked prefill, we switch to a decoding-optimized block graph while prefill computation continues in parallel. The switched decoding graph is invoked multiple times within the prefill stage to ensure result correctness. This design rests on two observations: (i) executing short-input graphs multiple times is equivalent to a single long-input execution, and (ii) graph loading can be overlapped with one-round chunked prefill computation. Consequently, sd.npu can transition to the decoding graph without additional load overhead, yielding a negligible graph-switching cost.
$\bullet$ In-Context Distribution Calibration (Section 3.3). Motivated by the observation that, during prefill, the model computes next-token logits for each context token that already align with its output distribution, sd.npu uses these logits to calibrate the task context by constructing a model-distributed token tree via depth-first search. In the subsequent retrieval-based drafting stage, sd.npu retrieves from this calibrated tree to mitigate lexical divergence between the context and the model’s expectations. Compared with costly precomputation or task-specific fine-tuning, this lightweight procedure incurs negligible overhead.
$\bullet$ NPU-Optimized Draft Reuse (Section 3.4). sd.npu extends draft construction by selectively reusing plausible tokens from previously rejected drafts based on a confidence-based strategy. This is motivated by the observation that rejected tokens may still capture correct semantics and be accepted in later decoding steps. By reusing such tokens, sd.npu increases effective draft length with useful content, improving NPU utilization and reducing verification costs.
@@
$\bullet$ We design and implement sd.npu, an efficient NPU-optimized on-device context-augmented generation framework with three novel techniques, bridging the gap between retrieval-based speculative decoding and NPU offloading.
$\bullet$ We conduct comprehensive evaluations across four datasets, three smartphones, and three LLMs, demonstrating that sd.npu delivers consistent and significant performance gains in diverse mobile scenarios.
@@ 3.1. Overview
Figure 7. Overview of sd.npu framework.:
Design goal. sd.npu accelerates context-augmented generation by NPU-optimized retrieval-based speculative decoding. This entails (1) deploying and scheduling two compute graphs optimized for prefill and decoding, (2) aligning task context with LLM distribution, and (3) enriching drafts with sufficient high-value tokens to fully utilize NPU resources.
Workflow. Figure 7 illustrates the system architecture and workflow of sd.npu.
  • Offline stage. sd.npu partitions the LLM into multiple blocks, each containing several transformer layers. It profiles the load and compute latency of each block through warm-up to determine scheduling strategy.
  • Online stage. Upon receiving a request, sd.npu fetches task-relevant database and performs prefill using a SOTA chunk-based approach (Xu et al., 2025b; Chen et al., 2025a), splitting the input into chunks processed by a long-sequence optimized graph. During prefill, progressive graph scheduling (Section 3.2) incrementally switches the prefill-optimized graph to the decoding-optimized graph with load overlapped by computation. After prefill, output logits are used for in-context distribution calibration (Section 3.3), augmenting the database with calibrated contexts. sd.npu then performs retrieval-based speculative decoding using the decoding-optimized graph and the calibrated database. It iteratively constructs a drafting tree and verifies it via a confidence-based strategy to identify and reuse valuable tokens (Section 3.4). After an end-of-sequence token is predicted, the final response is returned and stored in the database manager as a historical context for future requests.
Key techniques. sd.npu incorporates three techniques to align retrieval-based speculative decoding with NPU offloading:
@@ 3.2. Progressive Graph Scheduling
Switching graph progressively. sd.npu gradually switches $G^{1}$ to $G^{2}$ during prefill. The design builds on two insights:
  • Graph equivalence by repetition: executing $G^{2}$ multiple times on shorter inputs yields the same result as one execution of $G^{1}$ on longer inputs, allowing early switching. As sd.npu maximizes draft lengths (Section 3.4), $G^{2}$ will not be too short (e.g., chunk/2 or chunk/4 of $G^{1}$), which will not incur much performance degradation.
@@
Figure 8 illustrates progressive scheduling with $N=6$ and two input chunks. We suppose that sd.npu decides to overlap loading a block by two blocks of computation based on the profiling results. Initially, all blocks are loaded with $G^{1}$. sd.npu computes $\operatorname{P}_{1},\operatorname{P}_{2}$ with $G^{1}$ while loading $G^{2}$ for $\operatorname{P}_{6}$. Then $\operatorname{P}_{3},\operatorname{P}_{4}$ execute while $G^{2}$ loads for $\operatorname{P}_{5}$. To compute $\operatorname{P}_{5},\operatorname{P}_{6}$ under $G^{2}$, chunk 1 is split into sub-chunks matching $G^{2}$’s shape, with $G^{2}$ executed repeatedly. Prefill of chunk 1 finishes with output logits while half of the model is switched. The same procedure applies to chunk 2. In contrast, synchronous loading waits for all compute to finish, and naive asynchronous loading only overlaps in chunk 2, both incurring higher latency (Figure 8(b)).
Problem Formulation. Formally, given $N$ model blocks that sequentially execute prefill for $C$ chunks, sd.npu aims to determine the time to load each block, i.e., choosing blocks $\{\operatorname{P_{i},\dots}\}$ to overlap load overhead of $\operatorname{P_{j}}$ by their computations at each compute-load round. The objective is to minimize the overall latency, which is equal to the maximum of the prefill completion time and the switch completion time.
Scheduling methodology. This scheduling problem is NP-hard as it is essentially equivalent to the 2-Partition problem (Korf, 1998) (proofed in Appendix A). sd.npu addresses this via a greedy algorithm, choosing the least $k$ consecutive blocks $\{\operatorname{P}_{i},\dots,\operatorname{P}_{i+k-1}\}$ that overlap a block load to execute. This strategy enables fully masking the load overhead by computation to minimize switching overhead. sd.npu finds the smallest $k$ by estimating $\widehat{T}_{\text{load}}(\operatorname{P}_{j};G^{2})$ and $\widehat{T}_{\text{compute}}(\operatorname{P}_{i};G^{x}_{i})$ based on the profiling results. To hide load latency, the sum of the next $k$ blocks’ compute time must be larger than the load overhead of the last block not switched, formulated as:
@@
sd.npu greedily chooses the minimal feasible $k$ that satisfies Equation 1:
@@
Then sd.npu executes $\operatorname{P}_{i},\dots,\operatorname{P}_{i+k^{\star}-1}$ under current graphs while loading $G^{2}$ for $\operatorname{P}_{j}$. This process repeats until all blocks are switched.
@@ 3.3. In-Context Distribution Calibration
Workflow. After prefill, sd.npu first collects predicted logits for all positions. Then sd.npu builds a calibrated token tree representing the model distribution by depth-first search, which includes two steps:
@@ 3.4. NPU-Optimized Draft Reuse
Reuse plausible drafts. As highlighted in Section 2.1, the model output usually shares similar semantics with the task context despite existing lexical gaps. Although these gaps cause rejection due to token-level misalignment, the rejected tokens include correct semantics and may be accepted in the next verification. sd.npu identifies and reuses these plausible tokens to construct NPU-friendly drafts lengthened by task-relevant information.
@@
Reuse methodology. To leverage reusable tokens in the first two cases, sd.npu aims to: (1) determine the start and end positions $\gamma,\epsilon$ of the maximum reusable segment; (2) estimate the reuse lifetime $\delta$, i.e., how many iterations the reused segment should be reserved for verification.
Enumerating all possible reuse strategies is of $O(N^{3})$ complexity, which is not suitable for online scheduling. Therefore, sd.npu adopts a confidence-based and length-first reuse strategy, assuming that drafts aligned with model predictions are more likely to be accepted in future steps. Therefore, we identify and retain the longest segment of $\mathbf{x}$ such that $x_{i}=y_{i}$ for all $i\in[e+\gamma,e+\epsilon]$, by solving:
@@
By maximizing draft length with potentially acceptable tokens, this reuse mechanism directly enhances NPU utilization and reduces verification times. The overhead of sd.npu’s drafting and reusing processes is near ignorable with ¡3 ms per decoding step, as discussed in Section 4.4.
@@ 4.1. Implementation and Setup
Implementation. We implement sd.npu  on top of the mllm framework (Xu et al., 2025b) with approximately 6,000 lines of C/C++ code. We reuse the NPU backend of mllm, including neural kernels and memory managers, to perform prefill and draft verification. sd.npu inherits kernel-level and layer-level optimizations from mllm and further extends the system with speculative decoding capabilities. Our modular design allows easy porting to other mobile LLM inference frameworks that face NPU inefficiencies during decoding. To enable NPU-optimized speculative decoding, we augment mllm with customized kernels and sampling strategies. In order to accelerate drafting, we implement a suffix automaton to identify potential suffix matches (Mohri et al., 2009). To support NPU graph switching, we integrate the graph saving and loading APIs provided by the Qualcomm QNN SDK (Qualcomm, 2025a).
Hardware. We evaluate sd.npu on three representative smartphones with Qualcomm SoCs: Redmi K60 Pro (Snapdragon 8 Gen 2, Android 13, 12 GB RAM), Redmi K70 Pro (Snapdragon 8 Gen 3, Android 14, 16 GB RAM), and OnePlus 13 (Snapdragon 8 Gen Elite, Android 14, 24 GB RAM). All devices are evaluated under real mobile conditions with the CPU frequency governed by the Android OS’s dynamic voltage and frequency scaling controller.
@@
Baselines. We compare sd.npu against six baselines, all implemented atop mllm to eliminate performance gaps from backend design differences  [^1] Other existing mobile LLM inference frameworks with NPU offloading support such as PowerInfer-v2 (Xue et al., 2024) and HeteroLLM (Chen et al., 2025a) are not open-source, making them unsuitable for direct implementation. sd.npu, as a decoding optimization, is orthogonal and integrable with PowerInfer-v2 and HeteroLLM, so we do not compare it directly. EdgeLLM (Xu et al., 2025a) optimizes model-based SD mainly to alleviate memory constraints. With sufficient memory, its performance reduces to that of EAGLE and does not address the memory-bound issue of decoding, thus making it an uninformative baseline for our settings. . All baselines use NPU for prefill, which is a widely-adopted approach for both industry and academic (Xu et al., 2025b; Chen et al., 2025a; Li et al., 2024a; Xue et al., 2024). We categorize these baselines by whether they use CPU for decoding or apply speculative decoding:
  • (1)NPU vanilla: use NPU for both prefill and decoding222mllm (Xu et al., 2025b) and other on-device LLM engines  (Chen et al., 2025a; Xue et al., 2024) offload precision-sensitive operations, such as attention, to CPU/GPU to improve response quality. sd.npu follows this pattern even in the NPU vanilla mode.;
@@
Metrics and Configuration. We evaluate average per-token latency, energy consumption and peak memory. We exclude accuracy comparison as sd.npu achieves lossless generation (Leviathan et al., 2023a). Energy is measured via Android’s virtual file system under /sys/class/power_supply by profiling every 100ms. The Redmi K70 Pro is excluded from energy results as the device lacks root access. Experiments are repeated three times and we report the average numbers.